# Extracting Participation in Collective Action from Social Media

### v1

@@ A.3 Annotation Codebook
We are annotating excerpts from social media comments and looking for instances of participation in collective action, expressed at different levels. In some cases, background knowledge (e.g. veganism being a social movement that advocates for the end of animal suffering and/or limiting resources for animal products to contrast climate change) might be useful for the annotations.
We define a collective action problem as:
  • Human-generated issue. An issue caused by social, political or economic factors faced by group of people (or an individual as an example of a symptom within a broader issue). The issue of electing a political candidate is considered a human-generated issue. Note that such definition can be slightly stretched to include animal rights. The problem isn’t a natural disaster or accident but is the result of human actions, decisions or inaction (e.g. flooding is not a collective action problem, lack of government support to flooding is). Examples:
    • Collective action problem: “All of the depression treatment was basically just ’it’s all in your head.’ They would just force you to do shit and call it treatment. If I did not participate in a module they convinced me that I was fucking up and prolonging my treatment.” The message reports an individual example of a pattern of a systemic problem, making it a collective action problem. Note that the possibility of extending the specific example to a broader problem should be present (in this case “they would just force you to do shit” helps generalizing the issue).
    • Not a collective action problem: “You are actively harmful to people because of your refusal to listen to them or believe that their suffering could be worse than yours. You say you don’t believe that delusions don’t exist in response to me talking about self-harm that happened due to delusion. You suggest that I go off the medication that makes it so I don’t kill myself to try to make my own.” This focuses on an interpersonal conflict with one person invalidating the other’s mental health experience, without the link to a systemic collective issue.
    • Not a collective action problem: “Don’t want that to happen to you? Don’t drink and drive. And if you do, then submit to the tests as you admitted to do when you obtained a drivers license.” This message does not identify any larger systemic issue that would require collective action. It refers to “that” without specifying what the issue is, and it does not explicitly specify how drinking and driving would impact a broader group.
  • Current. The problem is not historical or set in the past or, if it is, it should still be ongoing or there should be a clear identification of how it has impact on the present or on current or future actions that can be taken to mitigate it (e.g. Vietnam war faults is not a collective action problem, the discrimination of black people that started with slavery in the US is).
  • With opportunity for collective action. The issue can potentially be mitigated through group action (e.g. public advocacy, policy change, collective efforts, protests, social mobilization) or collective individual action (e.g. recycling).
  • Involves shared responsibility. Individuals in the in-group or even out-group members can take some form of action that contributes to alleviating the impacts of the issue.
The message needs to clearly refer to a collective action problem, i.e. coders should be able to identify and report the issue from the message itself, or determine the collective action problem as a broader expansion of a symptom described in the message. If the messages does not clearly report a collective action problem but refers to a social movement that is commonly known to be related to collective action problems (e.g. veganism) or describes the author as an activist or willing to become an activist, given the semantic charge of the concepts the message is still valid as defined within a collective action problem umbrella. In any other case, the message should be classified as None.
Coders should be particularly careful when the message contains parts of discourse simply reported from others: if the author adds information on top of the reported speech that addition might change the direction of the reported message.
Note that more than one commitment level might apply. When in doubt, always choose the higher ranked level (i.e. if “Intention” and “Execution” both apply, choose “Execution”).
Problem-solution
The comment highlights the collective action problem described within the message as actually problematic, potentially blaming those responsibilities and/or suggesting a way to fix it. Note that the suggestion of solutions could be expressed both in terms of collective or individual actions or steps to be taken, and will be generally about outlining steps to be taken rather than motivating others to join a cause. Opinions themselves can help identifying an issue as problematic, but stances or opinions without a clear identification of a collective action problem as actually problematic do not count. A similar argument can be considered for simple reports of facts that mention a collective action problem without identifying the issue as problematic. e.g. “I support reproduction rights, and I am aware that everyone has their opinion on it.” discusses a collective action problem, that is reproduction rights, but it does not actually identify it as a problem in the message and thus it does not count as a “Problem-solution” message.
A message referring to a collective action problem can be seen as identifying the issue as problematic even through its tone, for instance if emotionally charged and critical. e.g. “I think there’s a couple of problems with what you are saying. One gender is shipped off to fight and die. The other stays home.” uses a quite strong language to talk about the collective action problem of gender roles. e.g. “Men work more than women, it’s a fact.” relates to the gender roles collective action problem but does not identify it as a problem.
There is a subtle difference between proposing solution and calling for action (see next class), but the main difference lies in the sense of urgency, need for immediate action and giving motivation for others to join a cause that the Call-to-action class expresses. Usually, message that use the imperative tense should be considered as leaning towards Call-to-action, and messages using “X should do Y” more towards Problem-solution, but the contextual difference can identify the message in one category or the other. e.g. “If you want to help with the discrimination issue, you should contact one of the organizations listed in the submission post.” is considered a proposed solution and thus part of the “Problem-solution” class. e.g. “Join one of the organizations listed in the submission post to help with the discrimination issue.” implies a sense of higher urgency and would be part of the “Call-to-action” class. While the semantic meaning of the message is the same as the one above, the framing makes the difference.
Call-to-action
The comment tries to convince the audience to take a specific action (advocating, organizing, protesting, boycotting, legal action, but also “standing strong” or “support”) in response to a perceived issue within a collective action problem or become activists and to engage others in the action itself by placing an emphasis on collective Participation in real-time efforts. The proposed action could be immediate or long-term and is framed as necessary for addressing a collective problem, issue, or injustice. This action is often presented as a moral or ethical imperative, described with a sense of urgency, and often calls for a collective involvement.
If the message contains a Call-to-action but there is no collective action problem or mention of activism/movement that could be solved or tackled with the action, for instance because there is simply an expression of stance or opinion by the commenter, then the message should not be coded as part of the Call-to-action class. e.g. “I think the Republicans are ruining the US. We need to stop them now.” expresses a sense of urgency and contains a Call-to-action but it does not refer to any collective action problem. Since we are not interested in extracting frames in general but specific participation in collective action that consider our definition of collective action problem, the message would not be coded as containing the Call-to-action class.
Intention
The commenter shares their own desire to do something or give help to mitigate a collective action problem or be involved as an activist. Note that the collective action problem and/or the reference of an activist initiative (or a political/social movement) should be clear. Otherwise, the message can be considered as Intentional in its framing but it is hard to distinguish it from mere Intention to act in a general context. e.g. “I’ll participate if I can choose my conditions.” is a message that is too generic and cannot be associated to any collective action problem to be solved or activist movement.
Not that conditional Intention still counts as part of the Intention class. e.g. “I’ll donate to the men’s rights movement if it is going to a cause that is worth it.”
Execution
The commenter is describing their personal experience taking tangible actions to mitigate a collective action problem or being an activist. Both online and offline actions apply. Note that simply stating that one has been “fighting”, “standing for”, “supporting” does not qualify as Execution because it could be simply interpreted as interest on the issue rather than tangible actions taken to solve it.
Execution expressed without a clear definition of a collective action problem, a movement or being an activist is not enough for a message to be classified in the Execution class. e.g. “I donated to them and would do it again.” it is not clear in which context we are talking about donations, and it might as well be that it is not related to a collective action problem at all. If a political or social movement hint would be included, then the message could be coded as containing the Execution class. Note that expressions such as “I am a climate change activist” count as part of the Execution class even if no action is clearly stated. Being an activist is charged enough for the expression to be considered Execution.
None
The comment does not fit into one of the previous categories; its purpose isn’t clear nor relevant to a collective action problem.
Table B2: Performance metrics for the multi-class classification task performed directly.
Centr. 0.37 0.13 0.04 0.05 0.38 0.2 0.36
BERT CS 0.36 0.51 0.07 0.09 0.41 0.29 0.40
Syn 0.4 0.49 0.17 0.28 0.62 0.39 0.56
ZS No def 0.35 0.51 0.04 0.36 0.38 0.33 0.37
Def 0.38 0.41 0.12 0.27 0.44 0.32 0.42
SFT (Def) CS 0.4 0.61 0.08 0.33 0.5 0.39 0.48
Ext. 0.34 0.41 0 0.1 0.49 0.27 0.45
Ext. + Syn I/E 0.34 0.34 0.11 0.33 0.51 0.33 0.46
CS + Syn I/E 0.39 0.66 0.2 0.47 0.55 0.45 0.52
CS + Syn A 0.39 0.56 0.17 0.48 0.56 0.43 0.52
DPO (Def) CS 0.41 0.48 0 0.4 0.63 0.39 0.57
Ext. 0.33 0.22 0.21 0 0.48 0.25 0.43
Ext. + Syn I/E 0.35 0.28 0.29 0.32 0.66 0.38 0.58
CS + Syn I/E 0.37 0.54 0.2 0.37 0.4 0.38 0.4
CS + Syn A 0.39 0 0.23 0.42 0.51 0.31 0.46
@@ Appendix B Results
B.1 Additional Experiments on Test Set
@@
Table B3: Performance metrics for the multi-class classification task for the LLM fine-tuning approaches. This table considers the multi-class task performed both directly and as a second step of the layered approach, without considering the definition of what a collective action problem is in the prompt.
@@
B.2 Additional Details on Topic Modeling

### v2

@@ A.3 Annotation Codebook
See annotation_codebook.pdf in the GitHub repo https://github.com/ariannap13/extract˙collective˙action.
Method Train. set Participation None Macro Avg. Weight. Avg.
P R P R P R P R
Dict. 0.27 0.67 0.77 0.38 0.52 0.53 0.64 0.46
Centr. 0.36 0.81 0.88 0.49 0.62 0.65 0.75 0.57
BERT CS 0.29 0.96 0.93 0.17 0.61 0.56 0.76 0.37
CS + Syn A 0.44 0.60 0.84 0.73 0.64 0.67 0.74 0.70
ZS no def 0.53 0.58 0.85 0.82 0.69 0.70 0.77 0.76
def 0.53 0.66 0.87 0.80 0.70 0.73 0.78 0.76
SFT CS (no def) 0.32 0.88 0.89 0.35 0.61 0.61 0.74 0.49
CS (def) 0.32 0.93 0.93 0.31 0.62 0.62 0.77 0.47
CS + Syn A (no def) 0.35 0.86 0.90 0.44 0.62 0.65 0.76 0.55
CS + Syn A (def) 0.36 0.90 0.93 0.43 0.64 0.67 0.78 0.55
DPO CS + Syn A (def) 0.31 0.95 0.94 0.27 0.62 0.61 0.77 0.44
Table B2: Precision (P) and recall (R) scores for the binary classification task.
@@ Appendix B Results
B.1 Additional Performance Metrics
Table B2 reports the metrics of precision and recall for the binary task, while Table B3 them for the multi-class task.
Method Train. Set Prob.-Sol. Call-to-A. Intention Execution None Macro Avg. Weigh. Avg.
P R P R P R P R P R P R P R
Centr. 0.33 0.45 0.19 0.13 0.03 0.09 0.08 0.15 0.84 0.73 0.29 0.31 0.69 0.63
BERT CS 0.30 0.47 0.52 0.62 0.13 0.09 0.67 0.15 0.84 0.73 0.49 0.41 0.72 0.66
CS + Syn A 0.30 0.43 0.63 0.62 0.12 0.27 0.27 0.31 0.84 0.73 0.43 0.47 0.71 0.66
ZS no def 0.31 0.38 0.61 0.49 0.06 0.18 0.46 0.46 0.82 0.75 0.45 0.45 0.70 0.66
def 0.32 0.49 0.83 0.26 0.08 0.09 0.44 0.31 0.85 0.78 0.50 0.38 0.73 0.68
SFT (def) CS 0.30 0.48 0.56 0.56 0.20 0.09 0.50 0.38 0.84 0.73 0.48 0.45 0.72 0.67
Ext 0.28 0.49 0.48 0.31 0.00 0.00 0.00 0.00 0.84 0.73 0.32 0.31 0.70 0.65
Ext + Syn I/E 0.29 0.47 0.44 0.36 0.22 0.18 0.43 0.23 0.84 0.73 0.44 0.39 0.71 0.65
CS + Syn I/E 0.30 0.45 0.54 0.56 0.15 0.18 0.45 0.38 0.84 0.73 0.46 0.46 0.71 0.66
CS + Syn A 0.31 0.46 0.64 0.64 0.24 0.36 0.55 0.46 0.84 0.73 0.51 0.53 0.72 0.67
DPO (def) CS 0.28 0.27 0.50 0.10 0.00 0.00 0.27 0.31 0.78 0.84 0.37 0.30 0.66 0.68
Ext 0.28 0.46 0.47 0.41 0.00 0.00 0.00 0.00 0.84 0.74 0.32 0.32 0.70 0.65
Ext + Syn I/E 0.27 0.45 0.35 0.21 0.00 0.00 0.33 0.08 0.82 0.74 0.36 0.29 0.68 0.64
CS + Syn I/E 0.31 0.46 0.72 0.54 0.18 0.27 0.44 0.62 0.84 0.74 0.50 0.52 0.72 0.67
CS + Syn A 0.33 0.45 0.50 0.69 0.30 0.27 0.42 0.62 0.84 0.74 0.48 0.55 0.72 0.68
Table B3: Precision (P) and recall (R) metrics for the multi-class classification task as a second step of the layered approach.
B.2 Additional Experiments on Test Set
@@
Centr. 0.37 0.13 0.04 0.05 0.38 0.2 0.36
BERT CS 0.36 0.51 0.07 0.09 0.41 0.29 0.40
CS + Syn A 0.4 0.49 0.17 0.28 0.62 0.39 0.56
ZS No def 0.35 0.51 0.04 0.36 0.38 0.33 0.37
Def 0.38 0.41 0.12 0.27 0.44 0.32 0.42
SFT (def) CS 0.4 0.61 0.08 0.33 0.5 0.39 0.48
Ext. 0.34 0.41 0 0.1 0.49 0.27 0.45
Ext. + Syn I/E 0.34 0.34 0.11 0.33 0.51 0.33 0.46
CS + Syn I/E 0.39 0.66 0.2 0.47 0.55 0.45 0.52
CS + Syn A 0.39 0.56 0.17 0.48 0.56 0.43 0.52
DPO (def) CS 0.41 0.48 0 0.4 0.63 0.39 0.57
Ext. 0.33 0.22 0.21 0 0.48 0.25 0.43
Ext. + Syn I/E 0.35 0.28 0.29 0.32 0.66 0.38 0.58
CS + Syn I/E 0.37 0.54 0.2 0.37 0.4 0.38 0.4
CS + Syn A 0.39 0 0.23 0.42 0.51 0.31 0.46
Table B4: F1 scores for the multi-class classification task performed directly.
@@
Table B5: F1 scores for the multi-class classification task for the LLM fine-tuning approaches. This table considers the multi-class task performed both directly and as a second step of the layered approach, without considering the definition of what a collective action problem is in the prompt.
B.3 Additional Details on Topic Modeling